OpenStudy (anonymous):

Given a 60-μC point charge located at origin. find total electric flux passing through: a) that portion of the sphere r=26 cm bounded by and b) the close surface defined by ρ = 26 cm and z = ±26 cm c) the plane z = 26 cm

OpenStudy (anonymous):

for my first part it like that as we all know that flux density is the charge per unt area here charge is 60uc so divide 60uc by (4*pi*r2) we get D= 7.06*10^-5 c/m2. NOW given is portion with r=26cm theta= 0 to pi/2 and phi = 0 to pi/2 calculate required region area with formula =double integral(r^2sintheta dtheta dphi) we get =.106m^2 now multiply D*Required region we get 7.5uc M I RIGHT PLZ tell me

OpenStudy (anonymous):

but i am confused in b and c part how to calculate area of a plane z=26cm. pz reply as soon as possible

OpenStudy (anonymous):

plz send answer to my mail id= engg.akd@gmail.com